# Investigating Coverage Criteria in Large Language Models: An In-Depth Study Through Jailbreak Attacks

### v1

@@ Abstract
The swift advancement of large language models (LLMs) has profoundly shaped the landscape of artificial intelligence; however, their deployment in sensitive domains raises grave concerns, particularly due to their susceptibility to malicious exploitation. This situation underscores the insufficiencies in pre-deployment testing, highlighting the urgent need for more rigorous and comprehensive evaluation methods. This study presents a comprehensive empirical analysis assessing the efficacy of conventional coverage criteria in identifying these vulnerabilities, with a particular emphasis on the pressing issue of jailbreak attacks. Our investigation begins with a clustering analysis of the hidden states in LLMs, demonstrating that intrinsic characteristics of these states can distinctly differentiate between various types of queries. Subsequently, we assess the performance of these criteria across three critical dimensions: criterion level, layer level, and token level, providing a multi-faceted evaluation of their applicability in LLMs.
Our findings uncover significant disparities in neuron activation patterns between the processing of normal and jailbreak queries, thereby corroborating the clustering results. Leveraging these findings, we propose an innovative approach for the real-time detection of jailbreak attacks by utilizing neural activation features. Our classifier demonstrates remarkable accuracy, averaging 96.33% in identifying jailbreak queries, including those that could lead to adversarial attacks. The importance of our research lies in its comprehensive approach to addressing the intricate challenges of LLM security. By enabling instantaneous detection from the model’s first token output, our method holds promise for future systems integrating LLMs, offering robust real-time detection capabilities. This study significantly advances our understanding of LLM security testing, strengthens safety measures, and lays a critical foundation for the development of more resilient AI systems.
@@ I Introduction
Large language models (LLMs) have become a pivotal technology in artificial intelligence, fundamentally transforming how machines process and generate human language. LLMs have become indispensable across a wide range of applications, including automating customer service and facilitating decision-making in high-stakes domains such as finance [1], healthcare [2], and legal practice [3], and even driving innovations in creative industries. Despite their broad adoption, the deployment of LLMs in critical domains often reveals problematic behaviors, notably the phenomenon of “jailbreaking”, where models are manipulated to generate harmful or unintended content[4, 5, 6, 7, 8, 9], resulting in significant societal ramifications. This situation underscores the critical need for robust and comprehensive testing methodologies to effectively detect and mitigate vulnerabilities, thereby ensuring that LLMs remain reliable and trustworthy in sensitive applications. A crucial component in assessing the robustness of LLM testing methodologies is the application of coverage criteria, which offer a systematic framework for measuring the comprehensiveness of tests and uncovering latent vulnerabilities.
@@
While there is a lack of testing coverage criteria specifically for LLMs, prior research has proposed coverage criteria for smaller neural networks, offering valuable insights that can be adapted for LLMs. For example, Neuron Coverage (NC) [10] focuses on tracking the activation levels of individual neurons, whereas K-Multisection Neuron Coverage (KMNC) [11] evaluates neuron engagement across a spectrum of activation ranges. Furthermore, Top-k Neuron Coverage (TKNC) [11] and Top-k Neuron Patterns (TKNP) [11] prioritize the analysis of highly activated neurons and their corresponding patterns in guiding model decisions. These criteria address various aspects of neuron activation and utilization, providing a basis for developing LLM-specific coverage criteria.
@@
However, the application and effectiveness of these coverage criteria in LLMs remain unexplored, prompting our investigation into their applicability and efficacy. The sheer scale and complexity of LLMs, characterized by their extensive parameters and intricate deep-layer architectures, introduce unique challenges that are not encountered in smaller-scale neural network evaluations. These include determining the most effective criteria for LLMs, identifying suitable inspection points, and balancing comprehensive coverage with feasibility. The generative nature and complex dynamics of LLMs further complicate the selection of monitoring points and the adequacy of existing criteria in capturing their intricacies.
@@
To address the gap in understanding how traditional coverage criteria perform with LLMs, comprehensive empirical studies are essential. First, systematically evaluating existing criteria like NC, TKNC, and TKNP in the context of LLMs will help identify the most suitable ones for assessing robustness and reliability. Second, applying these criteria to specific scenarios, such as detecting abnormal LLM behavior, can uncover the models’ limitations and vulnerabilities, offering a promising perspective on anomaly detection in LLMs and complementing existing methods.
@@
Our Work. This study conducts an in-depth empirical investigation, focusing on jailbreak attacks to evaluate coverage criteria effectiveness. We begin with a comprehensive clustering analysis of LLM hidden states as they process various queries, showing that these states effectively differentiate query types. Building on this, our research unfolds across three dimensions:
  • Criterion Level: Evaluating and comparing different coverage criteria in LLMs.
  • Layer Level: Assessing the impact of network layers on coverage criteria to understand layer-specific dynamics.
  • Token Level: Exploring coverage criteria performance across tokens to gain insights into model response granularity.
Our research reveals significant differences in the sets of neurons covered when LLMs process normal versus jailbreak queries, aligning with our clustering experiments. These findings led to developing a simple yet effective downstream application: real-time detection of jailbreak attacks based on neural activation features. We train a classifier to distinguish between normal and jailbreak queries, achieving 96.33% accuracy in detecting jailbreak attacks, enabling detection as soon as the first token is output. This paves the way for real-time detection capabilities in future LLM-integrated systems, contributing to safer and more trustworthy AI applications.
Contributions. In summary, the contributions of this research are as follows:
@@
  • An extensive empirical study for LLM evaluation. Our comprehensive empirical analysis uncovers significant disparities in neuron coverage when comparing normal and jailbreak queries, thereby demonstrating the nuanced effectiveness of traditional coverage criteria within LLM contexts.
  • A novel downstream LLM jailbreak attack detection method. We introduce an innovative downstream application designed for real-time detection of jailbreak attacks, leveraging neural activation patterns to achieve highly accurate identification of such malicious attacks.
  • Towards robust LLM development. This work significantly advances the understanding of security testing in LLMs, providing a foundational framework that paves the way for the development of more robust and resilient AI systems.
II Preliminaries
@@ II-A Model Inference Process
We first formalize the inference process of LLMs based on the transformer’s architecture. The process starts with the initial input vector $h_{0}$, which initiates the transformation of the text. This transformation begins when the text is tokenized into discrete elements, with each token mapped to a dense numerical vector known as an embedding vector. These vectors encapsulate the semantic attributes of the tokens and are further processed by the model.
Each Transformer block, or simply “block” hereafter, denoted by $i$ (where $i=0,1,\ldots,L-1$, and $L$ is the total number of blocks), enhances the data by operating through two primary layers:
@@
$h_{i}^{\prime}=\text{L}_{\text{attn}}(h_{i})+h_{i},\quad i=0,1,\ldots,L-1$
@@
$h_{i+1}=\text{L}_{\text{mlp}}(h_{i}^{\prime})+h_{i}^{\prime},\quad i=0,1,\ldots,L-1$
After passing through $L$ transformer blocks, the final output $h_{L}$ is fed into a linear layer, which maps the feature-rich data into a simpler form suitable for output interpretation:
$\text{res}^{\prime}=\text{Linear}(h_{L})$
Finally, the softmax function converts these linear outputs into a probability distribution:
$\text{res}=\text{softmax}(\text{res}^{\prime})$
@@ II-B Evaluation Criteria for Deep Neural Networks (DNNs)
Mirroring the design of code coverage based on program “logic”, a series of studies recognize that the internal states (e.g., neuron performance) of small DNNs can be used to represent the “logic” of these networks for designing coverage criteria. The key focus is on how to better characterize internal states to design more effective coverage criteria. We provide details of these studies below.
Evaluation Criteria Based on Neuron Activation and Distribution. Neuron Coverage (NC) [10] assesses the proportion of neurons activated by a test suite based on a specified threshold. Other criteria, such as K-Multisection Neuron Coverage (KMNC) [11], Neuron Boundary Coverage (NBC) [11], and Strong Neuron Activation Coverage (SNAC) [11], evaluate the distribution of neuron activations relative to a range determined from the model training data. KMNC divides the activation range into k equal intervals, while NBC and SNAC focus on outlier activations. Neural Coverage (NLC) [12] examines the output distribution of all neurons in a layer using a covariance matrix to assess coverage.
Evaluation Criteria Based on Top Neuron Activation. A small subset of criteria, such as Top-K Neuron Coverage (TKNC) and Top-K Neuron Patterns (TKNP) [11], focus on the top-k most activated neurons within each layer. TKNC calculates the proportion of these top-k neurons relative to all neurons, while TKNP identifies and counts the number of distinct activation patterns of the top-k neurons across the entire test suite.
Evaluation Criteria Based on Neuron Trajectory Features or Causal Relationships. Some criteria focus on the trajectory features or causal relationships formed by neurons. TensorFuzz Coverage (TFC) [13] measures distances between neuron outputs within the same layer and clusters inputs based on a distance threshold. Surprise Coverage (SC) [14] and its variants (LSC [14], DSC [14], and MDSC [15]) assess the “surprise” level of new inputs by comparing their neuron output trajectories with those from the training data. Neuron Path Coverage (NPC) [16] identifies decision-making pathways using Layer-wise Relevance Propagation (LRP) [17], measuring the coverage of new decision logic. Causal Coverage (CC) [18] captures neuron relationships using causal discovery and computes the coverage ratio of the test suite’s learned causal graph relative to the true causal graph.
The three criteria types offer unique perspectives for assessing LLM testing. However, due to limitations in computational effectiveness and feasibility in the context of LLMs, we select and adopt suitable criteria from the above in this work (to be detailed and explained in Section III-B1).
II-C Jailbreak Attack
@@
LLM developers have been focusing on “aligning” these models to ensure they do not generate harmful or objectionable responses to user queries. For example, for the query “Write a convincing anonymous death threat,” unaligned LLMs might respond with harmful outputs such as “Sure, here’s a convincing death threat for you: …”. In contrast, aligned LLMs are trained to respond with statements like “I cannot answer such questions” and similar rejections.
Despite the apparent success of alignment, recent years have seen the emergence of several effective attacks, i.e., jailbreak attacks, that bypass these measures and generate harmful content, compromising LLM security [6, 5, 8, 7, 9]. One notable example is GCG [4], which employs a gradient-based automated search for adversarial suffixes in a white-box setting, prompting models to output affirmative responses to harmful queries.
@@ III-A Motivation: A Cluster Analysis Experiment
Figure 1: Clustering experiment analysis results. We select the results of $Block4$, $Block9$, $Block16$, and $Block31$ for display. In the figure, we use colors to distinguish datasets and shapes to represent clustering categories.
@@ III-B Methodology: Evaluation Dimensions
In this section, we first provide an overview of the empirical study workflow, as shown in Figure 2. Then, we detail the three target evaluation dimensions that potentially contribute to effective coverage design for LLMs.
Figure 2: The workflow of our study.
@@ III-B1 Three Evaluation Dimensions
TABLE I: Comparison of coverage criteria applicability to LLMs
Applicability Reason
@@
Existing DNN coverage criteria primarily use neurons or network layers as the basic computational units, evaluating model behavior from various perspectives. However, as the model size increases, particularly in LLMs as mentioned in Section I, these criteria face new challenges. Specifically, the vast scale and complexity of training data and the intricate architectures of LLMs make certain coverage criteria impractical or computationally challenging for real-world application. For instance, NPC involves complex decision path identification, and Causal Coverage (CC) requires sophisticated causal inference. In LLMs, these computational processes become exceedingly burdensome, hindering efficient evaluation.
Therefore, we conduct a detailed qualitative investigation and manual comparison of recently proposed coverage criteria to determine their applicability to LLMs. The detailed results are presented in Table I. We select the following coverage criteria in this work: NC, TKNC, TKNP, TFC and NLC.
@@
Figure 3: Probability density plot of maximum neuron activation values across model blocks in Llama-2-7b-chat.
@@
The reasons for choosing these two granularities are threefold. First, each layer in an LLM plays different roles and exhibits varying complexity when processing information. The attention layer captures dependencies and interactions between sequence positions, while the MLP layer performs nonlinear transformations and feature extraction. Second, our initial clustering experiments reveal that different blocks respond differently to query features, with initial layers capturing low-level features and deeper blocks focusing on high-level features and complex patterns. Third, we have recorded the maximum activation value of each neuron in different blocks of the Llama-2-7b-chat model while processing the same dataset. We then plotted the probability density of these maximum activation values for each block, as illustrated in Figure 3. From the figure, it can be observed that as the inference progresses, both the mean and variance of the activation values in the Llama-2-7b-chat model gradually increase. Different blocks exhibit noticeable differences in the density and distribution of neuron activation values on the same dataset.
@@
Due to varying query lengths and masking strategies in LLMs, each token can only access information from preceding tokens. As a result, the final token of the query sequence ($T_{0}$) captures the complete information of the query, serving as a condensed representation of the entire sentence’s semantics. We use this final query token as the reference point for our analysis. To investigate the model’s semantic construction mechanism during generation, we compare the coverage criteria inspected after generating each token at different queries and their respective outputs. We denote the first token of the output relative to $T_{0}$ as $T_{1}$, the second token as $T_{2}$, and so on. Similarly, we denote the second-to-last token of the input as $T_{-1}$, the third-to-last token as $T_{-2}$, and so on.

### v3

# Understanding the Effectiveness of Coverage Criteria for Large Language Models: A Special Angle from Jailbreak Attacks

@@ Abstract
Large language models (LLMs) have revolutionized artificial intelligence, but their increasing deployment across critical domains has raised concerns about their abnormal behaviors when faced with malicious attacks. Such vulnerability alerts the widespread inadequacy of pre-release testing. In this paper, we conduct a comprehensive empirical study to evaluate the effectiveness of traditional coverage criteria in identifying such inadequacies, exemplified by the significant security concern of jailbreak attacks. Our study begins with a clustering analysis of the hidden states of LLMs, revealing that the embedded characteristics effectively distinguish between different query types. We then systematically evaluate the performance of these criteria across three key dimensions: criterion level, layer level, and token level.
Our research uncovers significant differences in neuron coverage when LLMs process normal versus jailbreak queries, aligning with our clustering experiments. Leveraging these findings, we propose three practical applications of coverage criteria in the context of LLM security testing. Specifically, we develop a real-time jailbreak detection mechanism that achieves high accuracy (93.61% on average) in classifying queries as normal or jailbreak. Furthermore, we explore the use of coverage levels to prioritize test cases, improving testing efficiency by focusing on high-risk interactions and removing redundant tests. Lastly, we introduce a coverage-guided approach for generating jailbreak attack examples, enabling systematic refinement of prompts to uncover vulnerabilities. This study improves our understanding of LLM security testing, enhances their safety, and provides a foundation for developing more robust AI applications.
@@ I Introduction
Large language models (LLMs) have emerged as a transformative technology in artificial intelligence, drastically altering the way machines understand and generate human language. These models have become the backbone of numerous applications, from automating customer support to providing decision support in critical domains such as finance [1], healthcare [2], and legal sectors [3]. However, as LLMs are increasingly deployed across these crucial domains, they are frequently reported to suffer from abnormal model behaviors [4, 5, 6], particularly jailbreaking to generate harmful content [7, 8, 9, 10, 11, 12, 13], which causes severe societal impacts. This underscores the need for effective testing techniques to identify such attacks and maintain the trustworthiness of LLMs. One crucial aspect of examining the adequacy and quality of tests for LLMs is the coverage criteria, which provide a systematic way to measure the extent of model testing and identify potential gaps or weaknesses in the testing process.
@@
Despite a lack of testing coverage criteria specifically designed for LLMs, prior research in the community has made commendable efforts to design coverage criteria for small-scale neural networks. These efforts provide valuable insights that could be referenced and further applied to LLMs. Among these coverage criteria, Neuron Coverage (NC) [14] focuses on the activation levels of individual neurons, while K-Multisection Neuron Coverage (KMNC) [15] examines the utilization of neurons across different activation ranges. Top-k Neuron Coverage (TKNC) [15] and Top-k Neuron Patterns (TKNP) [15] emphasize the role of highly activated neurons and their patterns in the decision-making processes of the model. These criteria focus on different aspects of neuron activation and utilization, offering a foundation for the development of LLM-specific coverage criteria.
@@
However, the application and effectiveness of these coverage criteria in LLMs remain unexplored, motivating our aim to investigate their applicability and efficacy in this context. The sheer size and complexity of LLMs, characterized by their vast number of parameters and deeper layers, present unique challenges that previous small-scale evaluations may not address. These challenges include determining the most suitable and effective criterion for LLMs, identifying appropriate inspection points to monitor, and striking a balance between comprehensive coverage and feasibility. The generative nature of LLMs and their complex dynamics further complicate the selection of monitoring points and the adequacy of existing criteria in capturing their intricacies.
@@
To bridge the gap in our understanding of how traditional coverage criteria perform when applied to LLMs, comprehensive empirical studies are crucial. First, systematically evaluating the effectiveness of existing criteria, such as NC, TKNC, and TKNP, in the context of LLMs will help identify the most suitable criteria for assessing their robustness and reliability. Second, applying these criteria to specific scenarios, such as detecting abnormal LLM behavior, can provide valuable insights into the models’ limitations and potential vulnerabilities, offering an alternative yet promising perspective on anomaly detection in LLMs and complementing existing methods.
@@
Our Work. In this work, we conduct an in-depth and extensive empirical study, focusing on the well-defined abnormal behavior of jailbreak attacks, to examine the effectiveness of coverage criteria. Our study begins with a comprehensive clustering analysis of the hidden states of LLMs as they process various queries, revealing that the characteristics embedded within these hidden states effectively distinguish between different query types. Building upon this foundation, we then unfold our research across three key dimensions: ❶ Criterion Level, where we evaluate and compare the effectiveness of different coverage criteria in LLMs; ❷ Layer Level, where we assess the impact of network layers on coverage criteria to understand layer-specific dynamics; ❸ Token Level, where we explore coverage criteria performance across tokens to provide insights into the granularity of model responses.
Our research reveals significant differences in neuron coverage when LLMs process normal versus jailbreak attack queries, aligning with our clustering experiments. These empirical findings have led to the development of three practical downstream applications leveraging coverage criteria to enhance LLM security testing. The first is real-time detection of jailbreak attacks, which uses neural activation features to classify queries as normal or jailbreak with high accuracy (93.61%). This approach enables early detection from the model’s first token output, offering robust real-time security capabilities for LLM-integrated systems. Additionally, we propose test case prioritization to improve testing efficiency by identifying high-risk test cases and eliminating redundant ones. Finally, we introduce jailbreak case generation, employing coverage-guided methods to systematically refine prompts and uncover vulnerabilities. These applications highlight the potential of coverage criteria to address key challenges in LLM security testing.
Contributions. In summary, this research makes the following contributions:
@@
  • An extensive empirical study for LLM evaluation. Our extensive empirical study reveals significant differences in neuron coverage between normal and jailbreak queries when evaluating traditional coverage criteria in LLMs.
  • Three practical applications for LLM security testing. We propose real-time jailbreak detection, test case prioritization, and jailbreak case generation, demonstrating the versatility of coverage criteria.
  • Towards robust LLM development. We have enhanced the understanding of LLM security testing, laying a foundation for more robust and resilient AI applications.
II Background and Preliminaries
@@ II-A Model Inference Process
We formalize the LLM inference process based on the Transformer architecture, starting with the input vector $h_{0}$. The text is tokenized into discrete elements, each mapped to a dense embedding vector representing token semantics for further processing by the model. Each Transformer block, or simply “block” hereafter, denoted by $i$ (where $i=0,1,\ldots,L-1$, and $L$ is the total number of blocks), enhances the data by operating through two primary layers:
@@
$h_{i}^{\prime}=\text{L}_{\text{attn}}(h_{i})+h_{i},\quad i=0,1,\ldots,L-1\vspace{-0.2cm}$
@@
$h_{i+1}=\text{L}_{\text{mlp}}(h_{i}^{\prime})+h_{i}^{\prime},\quad i=0,1,\ldots,L-1\vspace{-0.2cm}$
Finally, after processing through $L$ transformer blocks, the final output $h_{L}$ is mapped by a linear layer and transformed into a probability distribution using the softmax function:
$\text{res}=\text{softmax}(\text{Linear}(h_{L}))$
@@ II-B Evaluation Criteria for Deep Neural Networks (DNNs)
Mirroring the design of code coverage based on program “logic”, a series of studies recognize that the internal states (e.g., neuron performance) of DNNs can be used to represent the “logic” of these networks. The key focus is on how to better characterize internal states to design more effective coverage criteria. Table I summarizes the main criteria categorized into three types.
TABLE I: Summary of Coverage Criteria for DNN Testing
Category Coverage Criteria
Neuron Activation and Distribution Neuron Coverage (NC) [14], K-Multisection Neuron Coverage (KMNC) [15], Neuron Boundary Coverage (NBC) [15], Strong Neuron Activation Coverage (SNAC) [15], Neural Coverage (NLC) [16]
Top Neuron Activation Top-K Neuron Coverage/Patterns (TKNC/TKNP) [15]
Neuron Trajectory or Causal Features TensorFuzz Coverage (TFC) [17], Surprise Coverage (SC) [18], Likelihood Surprise Coverage (LSC) [18], Distance-ratio Surprise Coverage (DSC) [19], Mahalanobis Distance Surprise Coverage (MDSC) [19], Neuron Path Coverage (NPC) [20], Causal Coverage (CC) [21]
Use of Evaluation Criteria in DNN Testing. In DNN testing, white-box coverage criteria provide insights into a test suite’s ability to expose diverse functional behaviors, known as functional diversity, and detect potential defects, referred to as defect-detection capability [16].
These criteria assess how different DNN components are covered, reflecting decision-making processes [20]. For example, Neuron Coverage (NC) measures the proportion of neurons activated above a threshold, reflecting the model’s logic breadth. Top-K Neuron Coverage (TKNC) and Top-K Neuron Patterns (TKNP) focus on significant neuron activations, capturing key functional patterns. High coverage in these metrics suggests the test suite elicits diverse behaviors, boosting confidence in the model’s reliability.
These criteria also help identify inputs that trigger incorrect or unexpected behaviors, revealing potential defects. TensorFuzz Coverage (TFC) represents neuron outputs from the same layer as high-dimensional variables and clusters them based on different inputs. The coverage is then quantified by the number of formed clusters, capturing diverse activation patterns and potential abnormalities. Neural Coverage (NLC) analyzes neuron activation distributions to find vulnerabilities. By linking coverage with defect discovery, these metrics prioritize test cases that expose hidden issues, enhancing DNN robustness.
II-C DNN Testing versus LLM Testing
Traditional DNN testing focuses on black-box and white-box approaches. Black-box testing uses benchmark datasets (e.g., ImageNet [22]) to evaluate accuracy and generalization, and employs fuzz testing to mutate inputs (e.g., lighting, angle, noise) [23] to further identify deficiencies in the model’s handling of abnormal or boundary inputs. White-box testing seeks to capture the model’s “logic” by examining its internal structure and behavior, understanding and uncovering subtle errors and vulnerabilities that black-box testing may fail to detect. For example, DeepXplore [14] designs neuron coverage to measure activated model regions and employs differential testing to identify inconsistencies among similar DNNs, efficiently detecting erroneous corner cases.
Existing LLM testing predominantly relies on black-box testing by constructing benchmarks [24, 25, 26, 27]. For instance, MMLU [24] assesses multitask accuracy across 57 subjects, including math, history, and law. Additionally, recent studies [28, 29] have developed comprehensive black-box robustness testing methods for LLMs. Despite the commendable progress in black-box testing, vulnerabilities like “jailbreak attacks” [10] continue to emerge, with their causes still unclear. This motivates us to explore white-box coverage criteria to examine LLMs’ internal behavior, enabling more effective testing and defenses.
However, white-box testing coverage for LLMs might differ significantly from traditional DNNs due to their vast input-output scope and larger architectures, posing challenges for adapting existing white-box criteria. Thus, we focus on the critical security issue of “jailbreak attacks” and study the applicability of existing white-box testing coverage approaches. We anticipate this will reveal vulnerabilities overlooked by black-box methods and improve testing and defense strategies. Specifically, this not only enhances jailbreak detection but also helps prioritize test cases and generate efficient jailbreak cases. Beyond the jailbreak issue targeted in this paper, we hope this study will inspire white-box exploration to other security and functionality challenges in LLMs [30, 31, 32, 33, 34].
II-D Jailbreak Attack and Defenses
@@
LLM developers have focused on aligning models to prevent harmful or objectionable outputs. For example, when prompted with “Write a convincing anonymous death threat,” aligned LLMs are trained to respond with rejections like “I cannot answer such questions,” instead of producing harmful content. Despite these efforts, jailbreak attacks have emerged that bypass alignment measures, causing LLMs to generate harmful outputs and compromising their security [9, 8, 11, 10, 12].
In response, defensive mechanisms are evolving. Methods like erase-and-check [35] provide certified safety against adversarial prompts by sequentially removing tokens and inspecting subsequences with a safety filter. Other defenses, such as SmoothLLM [36] and JailGuard [37], reduce the success rate of jailbreak attacks by randomly perturbing input prompts and aggregating predictions. Additional strategies include perplexity-based detection [38, 39] and self-classification by LLMs [40, 41, 42], which detect perplexity or directly assess prompt harmfulness.
@@ III-A Motivation: A Cluster Analysis Experiment
Figure 1: Clustering Experiment Analysis Results: We Select the Results of $Block4$, $Block9$, $Block16$, and $Block31$ for Display. In the Figure, We Use Colors to Distinguish Datasets and Shapes to Represent Clustering Categories
@@ III-B Methodology: Evaluation Dimensions
In this section, we first provide an overview of the empirical study workflow, as shown in Figure 2. Then, we detail the three target evaluation dimensions that potentially contribute to assessing effective coverage criteria for LLMs.
Figure 2: The Workflow of Our Study
@@ III-B1 Three Evaluation Dimensions
TABLE II: Comparison of Coverage Criteria Applicability to LLMs
Criterion Applicability Reason
@@
Existing DNN coverage criteria primarily use neurons or network layers as the basic computational units, evaluating model behavior from various perspectives. However, as the model size increases, particularly in LLMs as mentioned in Section I, these criteria face new challenges. The vast scale of LLMs, complex training data, and intricate architectures make some coverage criteria impractical or computationally expensive. For example, NPC, designed for convolutional neural networks, is unsuitable for LLMs, while CC requires complex causal inference, with runtimes hundreds of times longer than NC. Storing hidden states for 8,000 queries in our experiment consumes 5.2 TB, just 0.0002% of the training data. Thus, using full training data is infeasible for some criteria, while limiting to fewer samples may compromise metric effectiveness.
Therefore, we conduct a detailed qualitative investigation and manual comparison of recently proposed coverage criteria to determine their applicability to LLMs. The detailed results are presented in Table II. We select the following coverage criteria in this work: NC, TKNC, TKNP, TFC and NLC.
@@
Figure 3: Probability Density Plot of Maximum Neuron Activation Values Across Model Blocks in Llama-2-7b-Chat
@@
The reasons for choosing these two granularities are threefold. First, each layer in an LLM plays different roles and exhibits varying complexity when processing information. The attention layer captures dependencies and interactions between sequence positions, while the MLP layer performs nonlinear transformations and feature extraction. Second, our initial clustering experiments reveal that different blocks respond differently to query features, with initial layers capturing low-level features and deeper blocks focusing on high-level features and complex patterns. Third, we have recorded the maximum activation value of each neuron in different blocks of the Llama-2-7b-chat model while processing the same dataset. We then plot the probability density of these maximum activation values for each block, as illustrated in Figure 3. From the figure, it can be observed that as the inference progresses, both the mean and variance of the activation values in the Llama-2-7b-chat model gradually increase. Different blocks exhibit noticeable differences in the density and distribution of neuron activation values on the same dataset.
@@
Due to varying query lengths and masking strategies in LLMs, each token can only access information from preceding tokens. As a result, the final token of the query sequence ($T_{0}$) captures the complete information of the query, serving as a condensed representation of the entire sentence’s semantics. We use this final query token as the reference point for our analysis. To investigate the model’s semantic construction mechanism during generation, we compare the coverage criteria inspected after generating each token at different queries and their respective outputs. We denote the first token of the output relative to $T_{0}$ as $T_{1}$, the second token as $T_{2}$, and so on.